# Seemingly unrelated Bayesian additive regression trees for cost-effectiveness analyses in healthcare

### v1

@@ 7 Discussion
Figure 10: Simulation results for continuous outcomes for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=250$ and $d=2$.
@@
Figure 11: Simulation results for continuous outcomes for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=250$ and $d=3$.
Figure 12: Simulation results for continuous outcomes for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=2$.
Figure 13: Simulation results for continuous outcomes for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=3$.
Figure 14: Simulation results for continuous outcomes for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=250$ and $d=2$.
Figure 15: Simulation results for continuous outcomes for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=2$.
Figure 16: Simulation results for continuous outcomes for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=1000$ and $d=3$.
Figure 17: Simulation results for continuous outcomes for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=3$.
Figure 18: Simulation results for continuous outcomes for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=1000$ and $d=3$.
Figure 19: Simulation results for binary outcomes for Friedman #3 with $n_{\text{train}}=n_{\text{test}}=250$ and $d=2$.
Figure 20: Simulation results for binary outcomes for Friedman #3 with $n_{\text{train}}=n_{\text{test}}=250$ and $d=3$.
Figure 21: Simulation results for binary outcomes for Friedman #3 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=2$.
Figure 22: Simulation results for binary outcomes for Friedman #3 with $n_{\text{train}}=n_{\text{test}}=500$ and $d=3$.
@@
Table 6: RMSE and coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=250$ for continuous outcomes.
@@
suBART mvBART BayesSUR suBART mvBART BayesSUR
@@
$\rho_{12}$ $0.04$ $0.39$ $0.30$ $0.00$
$\rho_{13}$ $0.07$ $0.32$ $0.39$ $0.00$
$\rho_{23}$ $0.08$ $0.18$ $0.39$ $0.02$
Table 7: RMSE coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=500$ for continuous outcomes.
RMSE CI coverage
suBART mvBART BayesSUR suBART mvBART BayesSUR
$d=2$
$\rho_{12}$ $0.03$ $0.07$ $0.34$ $0.47$ $0.07$ $0.00$
$d=3$
$\rho_{12}$ $0.02$ $0.38$ $0.48$ $0.00$
$\rho_{13}$ $0.04$ $0.32$ $0.39$ $0.00$
$\rho_{23}$ $0.05$ $0.17$ $0.45$ $0.00$
Table 8: RMSE and coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=250$ for continuous outcomes.
RMSE CI coverage
suBART mvBART BayesSUR suBART mvBART BayesSUR
$d=2$
$\rho_{12}$ $0.05$ $0.258$ $0.34$ $0.33$ $0.00$ $0.00$
$d=3$
$\rho_{12}$ $0.04$ $0.39$ $0.30$ $0.00$
$\rho_{13}$ $0.07$ $0.32$ $0.39$ $0.00$
$\rho_{23}$ $0.08$ $0.18$ $0.39$ $0.02$
Table 9: RMSE and coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=500$ for continuous outcomes.
RMSE CI coverage
suBART mvBART BayesSUR suBART mvBART BayesSUR
$d=2$
$\rho_{12}$ $0.04$ $0.21$ $0.50$ $0.32$ $0.00$ $0.00$
$d=3$
$\rho_{12}$ $0.05$ $0.56$ $0.24$ $0.00$
$\rho_{13}$ $0.07$ $0.49$ $0.33$ $0.01$
$\rho_{23}$ $0.06$ $0.71$ $0.55$ $0.00$
Table 10: RMSE and coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=1000$ for continuous outcomes.
RMSE CI coverage
suBART mvBART BayesSUR suBART mvBART BayesSUR
$d=2$
$\rho_{12}$ $0.06$ $0.17$ $0.49$ $0.02$ $0.00$ $0.00$
$d=3$
$\rho_{12}$ $0.05$ $0.58$ $0.02$ $0.00$
$\rho_{13}$ $0.05$ $0.49$ $0.18$ $0.01$
$\rho_{23}$ $0.03$ $0.72$ $0.51$ $0.00$
Table 11: RMSE and coverage of a 50% CI for $\bar{{\rho}}_{jk}$ from the posterior samples for Friedman #3 with $n_{\text{train}}=n_{\text{test}}=250$ for binary outcomes.

### v3

@@ 5.1 Continuous response experiments
Figure 2: Simulation results for Friedman #1 under the suBART, BART, mvBART, and BayesSUR models (from left to right) with $n_{\text{train}}=n_{\text{test}}=1000$ and $d=2$.
In terms of uncertainty estimation, Figure 2 shows that all methods exhibit reasonable coverage rates for the prediction intervals when $d=2$, except for the first outcome where the coverage for both mvBART and BayesSUR is notably high. Figure 3 also shows satisfactory coverage for both suBART and mvBART, while the coverage under BayesSUR is only close to the nominal rate when $j=3$, which is again due to the linearity of this response.
@@
Figure 3: Simulation results for Friedman #1 under the suBART, BART, and BayesSUR models (from left to right) with $n_{\text{train}}=n_{\text{test}}=1000$ and $d=3$.
For proper uncertainty quantification, it is essential to correctly estimate the elements of the covariance matrix $\boldsymbol{\Sigma}$. Table 3 displays the RMSE and coverage rates of 50% credible intervals for the correlation parameters $\rho_{jk}$ for all $j\neq k$ provided by suBART, mvBART, and BayesSUR. Notably, correlation values for the BART model are not provided as it assumes independence (i.e., $\rho_{jk}=0\>\forall\>j\neq k$), and the mvBART estimates are provided only when $d=2$, due to the aforementioned limitations of the skewBART package. Standard deviation parameters $\sigma_{j}$ are also given where applicable. From the results, it is clear that suBART outperforms BART, mvBART, and BayesSUR in terms of coverage, demonstrating its superior ability to estimate correlation structures. The coverage values of zero for BayesSUR are particularly notable and suggest an inability to accurately estimate correlations when assuming linear regressions for non-linear responses. In terms of RMSE, suBART is superior to BART and BayesSUR and comparable to mvBART, albeit only in the $d=2$ setting.
@@
Table 3: RMSE and coverage of a 50% CI for $\sigma_{k}$ and ${\rho}_{jk}$ for Friedman #1 with $n_{\text{train}}=n_{\text{test}}=1000$.
@@
suBART BART mvBART BayesSUR suBART BART mvBART BayesSUR
@@
$\sigma_{1}$ $0.02$ $0.04$ $1.62$ $0.46$ $0.28$ $0.00$
$\sigma_{2}$ $0.07$ $0.11$ $4.15$ $0.34$ $0.11$ $0.00$
$\sigma_{3}$ $0.18$ $0.19$ $0.10$ $0.25$ $0.22$ $0.41$
$\rho_{12}$ $0.01$ $0.34$ $0.39$ $0.00$
$\rho_{13}$ $0.02$ $0.31$ $0.45$ $0.00$
$\rho_{23}$ $0.03$ $0.16$ $0.50$ $0.00$
Equivalent figures and tables summarising the results for the remaining scenarios, with $d=3$ and/or different sample sizes, yield the same conclusions as above and have been deferred to Appendix C for brevity.
5.2 Binary response experiments
The experiments for binary responses are aligned with those from Section 5.1, wherein different sample sizes of $n_{\text{train}}=n_{\text{test}}=\{250,500,1000\}$ are used for training and test samples respectively. The simulation of the latent variables $z^{(j)}$ is described by the system of equations below. Other than when $j=3$, the values of the latent variables are non-linear functions. Recall that correlated noise $(\varepsilon_{i}^{(1)},\ldots,\varepsilon_{i}^{(d)})^{\top}\sim\text{MVN}_{d}(\mathbf{0}_{d},\boldsymbol{\Sigma})$ is subsequently added to the $d$-dimensional latent variable and that the generating process for each binary response follows Equation (3) thereafter. As the generative model is invariant to the values of the variances (Chib and Greenberg, 1998), we set them to $1$ for convenience, but the true correlations are set to the same values as in Table 2.
Friedman #2:
$z_{i}^{(1)}=\sin(x_{i}^{(1)}x_{i}^{(2)}\pi)+{x_{i}^{(3)}}^{3}$
$z_{i}^{(2)}=-1+2x_{i}^{(1)}x_{i}^{(4)}+e^{x_{i}^{(5)}}$
$z_{i}^{(3)}=0.5(x_{i}^{(2)}+x_{i}^{(4)})+x_{i}^{(5)}$
The models evaluated for these experiments are the probit suBART and probit extensions of the standard BART and BayesSUR. The mvBART model is excluded, since the skewBART software does not offer a probit implementation of the model, for any dimensionality. We persist in evaluating settings with varying dimension $d=\{2,3\}$ with the $d=2$ setting again comprising the first two responses. The results for $d=3$, with $n_{\text{train}}=n_{\text{test}}=1000$, are summarised in Figure 4 and are consistent with the findings from Section 5.1. When logarithmic loss and ACC are considered as metrics for evaluating predictive performance, suBART either exhibits superior results or comparable averages. Both tree-based models outperform Bayesian SUR, with the exception of the linear third response in the $d=3$ setting, as per the continuous-outcome simulation experiments in Section 5.1.
Regarding calibration, it is evident that suBART and BART are satisfactory across all scenarios, showing coverage rates close to the nominal value. On the other hand, due to the inherent linearity of BayesSUR, its calibration performance is notably poorer, though the third linear response in Figure 4 is again an exception in this regard, as per Section 5.1. As per the Friedman #1 simulations in Section 5.1, the results for remaining sample sizes are deferred to Appendix C, as they lead to similar conclusions.
Figure 4: Simulation results Friedman #2 with $n_{\text{train}}=n_{\text{test}}=1000$ and $d=3$.
The results regarding the estimation of the correlation parameters are presented in Table 4, which includes the RMSE and CI coverage for the correlation parameters $\rho_{jk}$ associated with binary responses when $n_{\text{train}}=1000$. As BART assumes independent errors, it is excluded from these results. Overall, these results are consistent with Table 3 in clearly demonstrating superior performance for suBART in terms of estimating correlation structures.
Table 4: RMSE and coverage of a 50% CI for ${\rho}_{jk}$ for Friedman #2 with $n_{\text{train}}=n_{\text{test}}=1000$.